OpenStudy (anonymous):

blue line is Y=x/2 -1 and green line is Y=2x+1 where do the lines cross ?

OpenStudy (nice):

Y = Y x/2 -1 = 2x+1 2x-x/2=-1-1 3x/2 = -2 3x = -4 x = -4/3

OpenStudy (mathteacher1729):

The lines cross where the two equations are equal: \[\frac{x}{2}-1=2x+1\] Can you solve that equation? Also, have you tried graphing these lines with geogebra or your calculator?

OpenStudy (nice):

put that x = -4/3 in any equation to get y comp.

OpenStudy (anonymous):

Let. blue line=red line Y=Y x/2-1=2x+1 x/2=2x+2 x=4x+4 x-4x=4 -3x=4 x=-4/3\ Answer=-4/3

OpenStudy (mathteacher1729):

Here is the graph.

OpenStudy (mathteacher1729):

You can create that graph by going to http://www.wolframalpha.com and typing in "plot y = x/2 -1 and y = 2x+1" (without the quotes).

OpenStudy (anonymous):

thanks for the help !